1.Find the x- and y-intercepts (if any) for each line
2.Plot those intercepts, and graph the two lines
3.Apply elimination or substitution to find the coordinates of the point of intersection (if there are no solutions or infinite solutions, state this
5x - 2y = -1
x + 4y = 35
i have a clue how to start. i really dont want the answer as much as i want the procedure for each step. i have more of these and i can figure out one then im in business

Answer by rfer(12657)

5x-2y=-1
x+4y=35, x=35-4y
5(35-4y)-2y=-1
175-20y-2y=-1
-22y=-176
y=8
----------------
5x-2(8)=-1
5x=15
x=3
(3,8) intersect
-------------------
5x-2y=-1
5(0)-2y=-1
y=-1/-2
y=1/2
(0,1/2) y int
-------------------------
5x-2(0)=-1
5x=-1
x=-1/5
(-1/5,0) x int